# Entropy Scaling for Diffusion Coefficients in Mixtures

### v1

@@ II.2 Diffusion Coefficients in Mixtures
Fig. 4a shows the entropy scaling predictions in comparison to the experimental reference data for both self-diffusion coefficients of the system $n$-hexane + $n$-dodecane at two temperatures. Both components strongly differ in their molar mass. Both self-diffusion coefficients decrease with increasing mole fraction $x_{2}$ in a non-linear way, which is predicted well by the framework. The entropy scaling predictions of both self-diffusion coefficients and the experimental data are in excellent agreement.
@@
Fig. 4b shows the results for the Fickian diffusion coefficient in the system toluene + $n$-hexane $D_{12}$ at four temperatures from 293.15 K to 308.15 K as a function of the mole fraction at ambient pressure. The Fickian diffusion coefficient increases with increasing temperature. Again, the entropy scaling predictions are in excellent agreement with the experimental data.
@@
Fig. 5 shows the predictions for the Maxwell-Stefan diffusion coefficient phase diagrams of the system toluene + $n$-hexane at two pressures: $p=0.1\,{\rm MPa}$ (top) and $p=4\,{\rm MPa}$ (bottom). The diffusion coefficients of the coexisting phases can be obtained in a straightforward manner by the framework. Additional results (including supercritical, metastable, and unstable states) are presented in the Supplementary Information. At $p=0.1\,{\rm MPa}$, both components are subcritical (see inset in Fig. 5 top). The Maxwell-Stefan diffusion coefficient is shown for three isotherms (340 K, 360 K, 385 K). As shown in the inset (see Fig. 5 top), one isotherm is entirely in the liquid phase (340 K), one isotherm passes through the vapor-liquid equilibrium (360 K), and one isotherm is entirely in the gas phase (385 K). The Maxwell-Stefan diffusion coefficient in the gas phase is significantly larger than that in the liquid phase. In the liquid phase, the diffusion coefficient strongly depends on the composition. In the gas phase, the Maxwell-Stefan diffusion coefficient exhibits only a minor dependency on the composition. This is in line with the Chapman-Enskog theory[2] – which is inherently incorporated in our framework. At $p=4\,{\rm MPa}$ (cf. Fig. 5 bottom), $n$-hexane is supercritical and a critical point exists at about $T=547.18\,{\rm K}$ (see inset). Four isotherms are plotted: One subcritical isotherm ($T=540\,{\rm K}$), the critical isotherm ($T=547.18\,{\rm}K$), one isotherm passing through the vapor-liquid equilibrium ($T=570\,{\rm K}$), and one isotherm entirely in the gas/supercritical phase ($T=590\,{\rm K}$). The isotherm at $T=540\,{\rm K}$ exhibits a transition between the two pseudo-pure component limiting cases – from a liquid state at $x_{2}=0$ to a supercritical state at $x_{2}=1\;{\rm mol\,mol^{-1}}$. At the critical point, the diffusion coefficient exhibits a large gradient with respect to the composition. Due to the thermodynamic consistency established by the EOS, the framework correctly predicts the Fickian diffusion coefficient to be zero at critical points (see Supplementary Information).
@@
Figure 5: Maxwell-Stefan diffusion coefficient Ð$\textit{\DH}_{12}$ in the mixture toluene (1) + $n$-hexane (2) as a function of the mole fraction of $n$-hexane $x_{2}$ at $p=0.1\,{\rm MPa}$ (top) and $p=4\,{\rm MPa}$ (bottom). Black lines: Vapor-liquid equilibrium; colored lines: Isotherms. Insets show the $T-x$ vapor liquid-equilibrium phase diagram at the respective pressure.
@@
In the Supplementary Information, it is shown that also diffusion coefficients of metastable and unstable states as well as at spinodal states can be described by the framework (which is for example relevant for nucleation). This significantly exceeds the capabilities of all presently available diffusion coefficient models.
III Discussion
@@
In this work, a new methodology for predicting diffusion coefficients in mixtures in a consistent manner for all fluid states is proposed. This includes self-diffusion and mutual diffusion coefficients. The new approach combines several physical concepts, including the Rosenfeld scaling, the Chapman-Enskog theory, and a molecular-based EOS. The framework enables predictions for the mixture diffusion coefficients without any adjustable parameters based on the limiting cases of the pure components and pseudo-pure components (i.e. infinite dilution). Due to the coupling of entropy scaling and EOS, the framework can consistently describe diffusion coefficients in mixtures in different phases (gas, liquid, supercritical, metastable) – including coexisting phases such as vapor-liquid and liquid-liquid equilibria. Also predictions to regions in which no data are available are possible. The strong predictive capabilities are a result of the physical backbone of the framework. The fact that infinite dilution diffusion coefficients exhibit a monovariate scaling behavior was discovered here for the first time. It enables an efficient entropy scaling modeling of the limiting mixture diffusion coefficient cases based on very few data. Since the limiting case diffusion coefficients are monovariate functions with respect to the entropy, extrapolations beyond the temperature and pressure range where data are available are feasible. In the mixture, no general monovariate behavior retains. Nevertheless, the mixture entropy in combination with mixing and combination rules enables predictions of mixture diffusion coefficients, which was demonstrated here for model mixtures as well as real substance systems.
@@
The new entropy scaling framework requires an accurate and robust EOS model that describes mixture thermodynamic properties reliably, i.e. the phase equilibria, the entropy, the thermodynamic factor, and the second virial coefficient. If such a model is available, the entropy scaling framework proposed in this work is a powerful tool. For future work, the extension to multi-component systems would be interesting.
IV Methods
@@
The framework proposed in this work consists of two new elements: (1) Treating the infinite dilution self-diffusion coefficient as a pseudo pure component to obtain a monovariate scaling; (2) The application of mixing and combination rules for predicting the different diffusion coefficients in a mixture.
IV.1 Entropy Scaling of Infinite Dilution Diffusion Coefficients
@@
The infinite dilution diffusion coefficient $D_{i}^{\infty}(T,p)$ is treated as a pseudo-pure component property such that the corresponding scaled property $\widehat{D}_{i}^{\infty,\circ}(s_{\rm conf})$ exhibits a monovariate function with respect to the configurational entropy $s_{\rm conf}$. For testing this postulate, the ES model for pure component diffusion coefficients from Ref. 16 is adopted here. Thereby, the infinite dilution self-diffusion coefficient is transformed using a modified Rosenfeld scaling (marked by ) as
$D_{i}^{\infty,\circ}=D_{i}^{\infty}\frac{\rho_{\rm N}^{1/3}}{\sqrt{RT/M_{\rm CE}}}\left(\frac{-s_{\rm conf}}{R}\right)^{2/3},$ (5)
@@
where $M_{i}$ and $M_{j}$ are the molar masses of the pure components. For the zero-density limit, the scaled Chapman-Enskog infinite dilution diffusion coefficient $D_{{\rm CE},i}^{\infty,\circ}$ is computed using a Lennard-Jones kernel (see Supplementary Information for details). Therefore, the Lennard-Jones cross-interaction size and energy parameters $\sigma_{ij}$ and $\varepsilon_{ij}$ are calculated by the Lorentz-Berthelot combining rules here [35, 36]. Finally, the CE-scaled infinite dilution diffusion coefficient $\widehat{D}^{\infty,\circ}_{i}$ is described by a split between the low-density (LD) and the high-density (HD) region as
@@
where $\tilde{s}_{\rm conf}=-s_{\rm conf}/(Rm)$ is the reduced configurational entropy using the segment parameter $m$ of molecular-based EOS and $W$ is a smoothed step function. The CE-scaled infinite dilution diffusion coefficient $\widehat{D}^{\infty,\circ}_{i}(\tilde{s}_{\rm conf})$ is then described by a function with up to five adjustable parameters $\alpha_{k}$ (see Supplementary Information). In this work, one or two system-dependent parameters were used for the real substance systems. A detailed description of the entropy scaling framework for infinite dilution diffusion coefficients is given in the Supplementary Information. To test and validate this new scaling approach for infinite dilution diffusion coefficients, two model systems and a real substance system (all strongly non-ideal systems) were studied in this work using molecular dynamics simulations using the molecular simulation engine ms2 [37]. Simulations were conducted in the all fluid regions, i.e. liquid, gaseous, supercritical, and metastable states (details are given in the Supplementary Information).
IV.2 Predicting Diffusion Coefficients in Mixtures
For predicting the diffusion coefficients $D_{i}(x_{j})$, $D_{j}(x_{j})$, Ð$\textit{\DH}_{ij}(x_{j})$, and $D_{ij}(x_{j})$ in a mixture, the entropy scaling framework is extended using combining and mixing rules. We briefly outline how this is done here; details are given in the Supplementary Information. Diffusion coefficients in mixtures are predicted based on the limiting case models for the self-diffusion coefficient of pure components and the (infinite dilution) pseudo-pure component. The scaled diffusion coefficient $\widehat{\Lambda}^{\circ}\in\{\widehat{D}_{i}^{\circ},\widehat{D}_{j}^{\circ},\widehat{\textit{\DH}}_{ij}^{\circ}\}$ is computed using the configurational entropy of the mixture $s_{\rm conf}^{\rm mix}=s_{\rm conf}^{\rm mix}(T,\rho,\underline{x})$ and a mixture model for the scaled function $\widehat{\Lambda}^{\circ}=\widehat{\Lambda}^{\circ}(s_{\rm conf}^{\rm mix},\alpha_{k,{\rm mix}})$, which is computed from the limiting case scaling functions. The final (unscaled) diffusion coefficient $\Lambda\in\{D_{i},D_{j},\textit{\DH}_{ij}\}$ is calculated as
@@ IV.2 Predicting Diffusion Coefficients in Mixtures
$\Lambda=\frac{\widehat{\Lambda}^{\circ}}{\frac{W(\tilde{s}_{\rm conf}^{\rm mix})}{\Lambda^{\circ}_{\rm CE}}+\frac{1-W(\tilde{s}_{\rm conf}^{\rm mix})}{\min\left(\Lambda^{\circ}_{\rm CE}\right)}}\cdot\frac{\sqrt{RT/M_{\rm ref}}}{\rho_{\rm N}^{1/3}}\left(\frac{-s_{\rm conf}^{\rm mix}}{R}\right)^{-2/3},$ (8)
@@
where $\Lambda^{\circ}_{\rm CE}$ is the Chapman-Enskog property of the mixture, $\rho_{\rm N}$ is the number density of the mixture, and $M_{\rm ref}$ is the reference mass of the mixture. The Fickian diffusion coefficient is obtained from the Maxwell-Stefan diffusion coefficient by the thermodynamic factor as predicted by the EOS (see Eq. 1). Details on the calculation of the Chapman-Enskog property of the mixture $\Lambda^{\circ}_{\rm CE}$, the reference mass $M_{\rm ref}$, and $\widehat{\Lambda}^{\circ}=\widehat{\Lambda}^{\circ}(s_{\rm conf}^{\rm mix},\alpha_{k,{\rm mix}})$ are given in the Supplementary Information. All required quantities are obtained using predictive combination rules and mixing rules and the EOS mixture model. No adjustable mixture parameters are introduced. The non-ideality of the mixture is primarily taken into account by the underlying EOS via the predicted mixture entropy. The combination of the EOS and entropy scaling enables the consistent calculation of the diffusion coefficients ($D_{i}$, $D_{j}$, Ð$\textit{\DH}_{ij}$, $D_{ij}$), homogeneous bulk properties ($pvT$, $s_{\rm conf}$, $\Gamma_{ij}$, etc.), and phase equilibria (e.g. vapor-liquid and liquid-liquid equilibria).
Supplementary Information
Supplementary Information with additional results and technical details is available.

### v2

@@ Results
The entropy scaling predictions of both self-diffusion coefficients and the experimental data are in good agreement. Only for the system acetone + chloroform, some deviations are observed. For the systems $n$-hexane + $n$-dodecane and nitrobenzene + $n$-hexane, the model predicts the experimental data very well; the mean relative deviations are $\overline{\updelta D_{1}}=1.24\,\%$ and $\overline{\updelta D_{2}}=2.91\,\%$ ($n$-hexane + $n$-dodecane) and $\overline{\updelta D_{1}}=4.96\,\%$ (nitrobenzene + n-hexane), respectively. For the system $n$-hexane + $n$-dodecane, the two components strongly differ in their molar mass, which yields a non-linear behavior of $D_{1}(x_{2})$ and $D_{2}(x_{2})$. This behavior is predicted well by the framework. Also for the system nitrobenzene + hexane, the non-ideality of the diffusion coefficients is well captured by the model. For the system acetone + chloroform, the mean relative deviations are $\overline{\updelta D_{1}}=8.58\,\%$ and $\overline{\updelta D_{2}}=9.69\,\%$.
Figure 6: Fickian diffusion coefficients of real substance mixtures predicted by the entropy scaling model as a function of the mole fraction $x_{2}$ at $p=0.1\,{\rm MPa}$. Symbols are experimental data and lines are model predictions. a) Mixture toluene (1) + $n$-hexane (2); experimental data from Ref. [47]. b) Mixture acetone (1) + chloroform (2); experimental data from Refs. [48, 49]. c) Mixture toluene (1) + acetonitrile (2); experimental data from Ref. [50]. Source data are provided as a Source Data file.
@@
Fig. 6 shows the results for the Fickian diffusion coefficient $D_{12}$ in three systems, namely toluene + $n$-hexane, acetone + chloroform, and toluene + acetonitrile. All results are at ambient pressure, but at different temperatures and shown as a function of the mole fraction. The Fickian diffusion coefficient increases with increasing temperature. The entropy scaling predictions are in good agreement with the experimental data ($\overline{\updelta D_{12}}=3.02\,\%$ for toluene + $n$-hexane, $\overline{\updelta D_{12}}=3.67\,\%$ for acetone + chloroform, and $\overline{\updelta D_{12}}=10.27\,\%$ for toluene + acetonitrile). For the system toluene + $n$-hexane, the entropy scaling model yields good agreement. For the systems toluene + acetonitrile and acetone + chloroform, some deviations are observed - especially in the vicinity of the extremum values.
@@
Fig. 7 shows the predictions for the Maxwell-Stefan diffusion coefficient phase diagrams of the system toluene + $n$-hexane at two pressures: $p=0.1\,{\rm MPa}$ (a) and $p=4\,{\rm MPa}$ (b). The diffusion coefficients of the coexisting phases can be obtained in a straightforward manner by the framework. Additional results (including supercritical, metastable, and unstable states) are presented in Suppl. Note 10. At $p=0.1\,{\rm MPa}$, both components are subcritical (see inset in Fig. 7 a). The Maxwell-Stefan diffusion coefficient is shown for three isotherms (340, 360, 385 K). As shown in the inset (see Fig. 7 a), one isotherm is entirely in the liquid phase (340 K), one isotherm passes through the vapor-liquid equilibrium (360 K), and one isotherm is entirely in the gas phase (385 K). The Maxwell-Stefan diffusion coefficient in the gas phase is significantly larger than that in the liquid phase. In the liquid phase, the diffusion coefficient strongly depends on the composition. In the gas phase, the Maxwell-Stefan diffusion coefficient exhibits only a minor dependency on the composition. This is in line with the Chapman-Enskog theory[2] – which is inherently incorporated in our framework. At $p=4\,{\rm MPa}$ (cf. Fig. 7 b), $n$-hexane is supercritical and a critical point exists at about $T=547.18\,{\rm K}$ (see inset). Four isotherms are plotted: One subcritical isotherm ($T=540\,{\rm K}$), the critical isotherm ($T=547.18\,{\rm K}$), one isotherm passing through the vapor-liquid equilibrium ($T=570\,{\rm K}$), and one isotherm entirely in the gas/supercritical phase ($T=590\,{\rm K}$). The isotherm at $T=540\,{\rm K}$ exhibits a transition between the two pseudo-pure component limiting cases – from a liquid state at $x_{2}=0$ to a supercritical state at $x_{2}=1\;{\rm mol\,mol^{-1}}$. At the critical point, the diffusion coefficient exhibits a large gradient with respect to the composition. Due to the thermodynamic consistency established by the EOS, the framework correctly predicts the Fickian diffusion coefficient to be zero at critical points (see Fig. 4).
@@
Figure 7: Maxwell-Stefan diffusion coefficient Ð$\textit{\DH}_{12}$ in the mixture toluene (1) + $n$-hexane (2) as a function of the mole fraction of $n$-hexane $x_{2}$ at $p=0.1\,{\rm MPa}$ (a) and $p=4\,{\rm MPa}$ (b). Black lines: Vapor-liquid equilibrium; colored lines: Isotherms. Insets show the $T-x$ vapor liquid-equilibrium phase diagram at the respective pressure. Source data are provided as a Source Data file.
@@
In Suppl. Note 10, it is shown that also diffusion coefficients of metastable and unstable states as well as at spinodal states can be described by the framework (which is for example relevant for nucleation). This significantly exceeds the capabilities of all presently available diffusion coefficient models.
Discussion
@@
In this work, a methodology for predicting diffusion coefficients in mixtures in a consistent manner for all fluid states is proposed. This includes self-diffusion and mutual diffusion coefficients. The approach combines several physical concepts, including the modified Rosenfeld scaling, the Chapman-Enskog theory, and a molecular-based EOS. The framework enables predictions for the mixture diffusion coefficients without any adjustable parameters based on the limiting cases of the pure components and pseudo-pure components (i.e. infinite dilution). Due to the coupling of entropy scaling and an EOS, the framework can consistently describe diffusion coefficients in mixtures in different phases (gas, liquid, supercritical, metastable) – including coexisting phases such as vapor-liquid and liquid-liquid equilibria. Even predictions to regions in which no data are available are possible. The strong predictive capabilities are a result of the physical backbone of the framework. The fact that infinite-dilution diffusion coefficients exhibit a monovariate scaling behavior was uncovered here for the first time. It enables an efficient entropy scaling modeling of the limiting mixture diffusion coefficient cases based on very few data. Since the limiting case diffusion coefficients are monovariate functions with respect to the entropy, extrapolations beyond the temperature and pressure range where data are available are feasible. In the mixture, no general monovariate behavior retains. Nevertheless, the mixture entropy in combination with mixing and combination rules enables predictions of mixture diffusion coefficients, which was demonstrated here for model mixtures as well as real substance systems.
@@
The entropy scaling framework requires an accurate and robust EOS model that describes mixture thermodynamic properties reliably, i.e. the phase equilibria, the entropy, the thermodynamic factor, and the second virial coefficient. If such a model is available, the entropy scaling framework proposed in this work is a powerful tool. For future work, the extension to multi-component systems, e.g. Refs. [51, 52, 53], would be interesting.
Methods
@@
The framework proposed in this work consists of two elements: (1) Treating the infinite dilution self-diffusion coefficient as a pseudo-pure component to obtain a monovariate scaling; (2) The application of mixing and combination rules for predicting the different diffusion coefficients in a mixture. Throughout this work, the configurational entropy $s_{\rm conf}(T,\rho,\underline{x})$ is calculated from molecular-based EOS as the derivative of the configurational Helmholtz energy $a_{\rm conf}$ with respect to temperature $T$ at constant volume $v$ and composition $\underline{x}$, i.e.
$s_{\rm conf}=-\left(\frac{\partial a_{\rm conf}}{\partial T}\right)_{v,\underline{x}}.$ (5)
Therein, the configurational Helmholtz energy $a_{\rm conf}$ is defined as $a_{\rm conf}=a-a_{\rm ideal}$ with $a_{\rm ideal}$ being the Helmholtz energy of the ideal gas. We use a reduced configurational entropy defined as
$\tilde{s}_{\rm conf}=-s_{\rm conf}/({\rm R}m),$ (6)
where $m$ is the segment parameter of molecular-based EOS.
Entropy scaling of infinite-dilution diffusion coefficients
@@
The infinite-dilution diffusion coefficient $D_{i}^{\infty}(T,p)$ is treated as a pseudo-pure component property such that the corresponding scaled property $\widehat{D}_{i}^{\infty,\circ}(\tilde{s}_{\rm conf})$ exhibits a monovariate function with respect to the reduced configurational entropy $\tilde{s}_{\rm conf}$ (cf. Eq. (6)). The modeling methodology is described in the following. There are many scaling approaches available in the literature that aim at establishing a monovariate behavior of transport coefficients. Here, we use the scaling described in Ref. [17]. The pure component scaling described in Ref. [17] (which follows Refs. [13, 16]) is adapted here to model the pseudo-pure component diffusion coefficients at infinite dilution. Thus, we propose that the infinite-dilution diffusion coefficient is transformed using
$D_{i}^{\infty,\circ}=D_{i}^{\infty}\frac{\rho_{\rm N}^{1/3}}{\sqrt{{\rm R}T/M_{\rm CE}}}\left(\frac{-s_{\rm conf}}{{\rm R}}\right)^{2/3},$ (7)
@@
where $M_{i}$ and $M_{j}$ are the molar masses of the pure components. The modified Rosenfeld-scaled diffusion coefficient $D_{i}^{\infty,\circ}$ exhibits some scattering in the zero-density limit for $s_{\rm conf}\rightarrow 0$. Therefore, the scaling is further modified using the Chapman-Enskog diffusion coefficient $D_{{\rm CE},i}^{\infty,\circ}$, which is computed using a Lennard-Jones kernel (see Suppl. Note 1 for details). Finally, the CE-scaled infinite-dilution diffusion coefficient $\widehat{D}^{\infty,\circ}_{i}$ is described by a split between the low-density (LD) and the high-density (HD) region as
@@
where $W$ is a smoothed transition function given by $W(\tilde{s}_{\rm conf})=1/(1+\exp(20(\tilde{s}_{\rm conf}-\tilde{s}_{\rm conf}^{\times})))$ with $\tilde{s}_{\rm conf}^{\times}=0.5$. The function $W$ establishes a smooth transition from the low-density region to the high-density region. The choice of $\tilde{s}_{\rm conf}^{\times}$ is heuristic. The resulting CE-scaled infinite-dilution diffusion coefficient can be described by a simple continuous, monovariate function $\widehat{D}^{\infty,\circ}_{i}=\widehat{D}^{\infty,\circ}_{i}(\tilde{s}_{\rm conf})$. Here, a function with two adjustable, system-dependent parameters $\alpha_{2}^{(D^{\infty}_{i})}$ and $\alpha_{3}^{(D^{\infty}_{i})}$ was used given as
@@
$\ln\left(\widehat{D}^{\infty,\circ}_{i}\right)=\frac{\alpha_{2}^{(D^{\infty}_{i})}(\tilde{s}_{\rm conf})^{2}+\alpha_{3}^{(D^{\infty}_{i})}(\tilde{s}_{\rm conf})^{3}}{1+g_{1}^{(D)}\ln(\tilde{s}_{\rm conf}+1)+g_{2}^{(D)}\tilde{s}_{\rm conf}},$ (10)
where $g_{1}^{(D)}=0.6632$ and $g_{2}^{(D)}=9.4714$ are global parameters fitted to the self-diffusion coefficient of the Lennard-Jones fluid [17]. The system-dependent parameters $\alpha_{2}^{(D^{\infty}_{i})}$ and $\alpha_{3}^{(D^{\infty}_{i})}$ were fitted to reference data of the pseudo-pure component, i.e. the diffusion coefficients at infinite dilution. This scaling approach was tested using model systems and real substance systems (see Results). Therefore, MD simulations were carried out using the simulation engine ms2 [54].
Predicting Diffusion Coefficients in Mixtures
For predicting the diffusion coefficients $D_{i}(x_{j})$, $D_{j}(x_{j})$, Ð$\textit{\DH}_{ij}(x_{j})$, and $D_{ij}(x_{j})$ in a mixture, the entropy scaling framework is extended using combining and mixing rules. Diffusion coefficients in mixtures are predicted based on the limiting case models, i.e. the models for the self-diffusion coefficient of the pure components and the (infinite dilution) pseudo-pure component. The self-diffusion coefficient $D_{i}$ in a mixture is calculated using the limiting case entropy scaling model of the pure component self-diffusion coefficient $D_{i}^{\rm pure}$ and the limiting case entropy scaling model of the infinite-dilution diffusion coefficient $D_{i}^{\infty}$ in the solvent $j$. The Maxwell-Stefan diffusion coefficient is calculated using both infinite-dilution diffusion coefficient limiting case entropy scaling models $D_{i}^{\infty}$ and $D_{j}^{\infty}$. Therefore, the scaled mixture diffusion coefficient $\widehat{\Lambda}^{\circ}\in\{\widehat{D}_{i}^{\circ},\widehat{D}_{j}^{\circ},\widehat{\textit{\DH}}_{ij}^{\circ}\}$ is computed as
$\ln\left(\widehat{\Lambda}^{\circ}\right)=\frac{\alpha_{2}^{(\Lambda)}(\tilde{s}_{\rm conf})^{2}+\alpha_{3}^{(\Lambda)}(\tilde{s}_{\rm conf})^{3}}{1+g_{1}^{(D)}\ln(\tilde{s}_{\rm conf}+1)+g_{2}^{(D)}\tilde{s}_{\rm conf}},$ (11)
where $\tilde{s}_{\rm conf}$ is the scaled configurational entropy of the mixture, i.e. $\tilde{s}_{\rm conf}=s_{\rm conf}/({\rm R}m_{\rm mix})$ with $m_{\rm mix}=x_{i}m_{i}+x_{j}m_{j}$. The parameters $\alpha_{k}^{(\Lambda)}$ (with $k\in\{2,3\}$) are calculated using linear mixing rules
$\alpha_{k}^{(D_{i})}$ $=x_{i}\alpha_{k}^{(D_{i}^{\rm pure})}+x_{j}\alpha_{k}^{(D_{i}^{\infty})},$ (12)
$\alpha_{k}^{(D_{j})}$ $=x_{i}\alpha_{k}^{(D_{j}^{\infty})}+x_{j}\alpha_{k}^{(D_{j}^{\rm pure})},$ (13)
$\alpha_{k}^{(\textit{\DH}_{ij})}$ $=x_{i}\alpha_{k}^{(D_{j}^{\infty})}+x_{j}\alpha_{k}^{(D_{i}^{\infty})},$ (14)
for the self-diffusion coefficients of components $i$, the self-diffusion coefficient of component $j$, and the Maxwell-Stefan diffusion coefficient, respectively.
The final (unscaled) diffusion coefficient $\Lambda\in\{D_{i},D_{j},\textit{\DH}_{ij}\}$ is calculated as
$\Lambda=\frac{\widehat{\Lambda}^{\circ}}{\frac{W(\tilde{s}_{\rm conf})}{\Lambda^{\circ}_{\rm CE}}+\frac{1-W(\tilde{s}_{\rm conf})}{\min\left(\Lambda^{\circ}_{\rm CE}\right)}}\cdot\sqrt{\frac{{\rm R}T}{M_{\rm ref}^{\Lambda}}}\frac{1}{\rho_{\rm N}^{1/3}}\left(\frac{-s_{\rm conf}}{{\rm R}}\right)^{-2/3},$ (15)
where $\Lambda^{\circ}_{\rm CE}$ is the Chapman-Enskog diffusion coefficient of the mixture, $\rho_{\rm N}$ is the number density of the mixture, and $M_{\rm ref}^{\Lambda}$ is the reference mass of the mixture, which is calculated as
$M_{\rm ref}^{D_{i}}$ $=x_{i}M_{i}+x_{j}M_{\rm CE},$ (16)
$M_{\rm ref}^{D_{j}}$ $=x_{i}M_{\rm CE}+x_{j}M_{j},\quad{\rm and}$ (17)
$M_{\rm ref}^{\textit{\DH}_{ij}}$ $=M_{\rm CE}$ (18)
for the self-diffusion coefficients $D_{i}$ and $D_{j}$ and the Maxwell-Stefan diffusion coefficient Ð$\textit{\DH}_{ij}$, respectively. Details on the calculation of the Chapman-Enskog property of the mixture $\Lambda^{\circ}_{\rm CE}$, the reference mass $M_{\rm ref}$, and $\widehat{\Lambda}^{\circ}=\widehat{\Lambda}^{\circ}(s_{\rm conf}^{\rm mix},\alpha_{k,{\rm mix}})$ are given in Suppl. Note 1.
@@
The Fickian diffusion coefficient is computed from the Maxwell-Stefan diffusion coefficient by the thermodynamic factor as predicted by the EOS, see Eq. (1). Thus, all required quantities are obtained using predictive combination rules and mixing rules and the EOS mixture model. No adjustable mixture parameters are introduced. However, for modeling the diffusion coefficients of a given binary system based on a given EOS model, adjustable parameters have to be determined for the four limiting cases, i.e. the two pure components and the two infinite dilution pseudo-pure components, cf. $\alpha$ parameters in Eq. (10). For predicting the diffusion coefficients in the mixture, no adjustable parameters are required. The non-ideality of the mixture is primarily taken into account by the underlying EOS via the predicted configurational mixture entropy. The combination of the EOS and entropy scaling enables the consistent calculation of the diffusion coefficients ($D_{i}$, $D_{j}$, Ð$\textit{\DH}_{ij}$, $D_{ij}$), homogeneous bulk properties ($pvT$, $s_{\rm conf}$, $\varGamma_{ij}$, etc.), and phase equilibria (e.g. vapor-liquid and liquid-liquid equilibria).
Acknowledgements
Author Contribution Statement
S. Schmitt: Data curation, Formal analysis, Methodology, Software (lead), Visualization, Writing – original draft. H. H.: Funding acquisition, Writing – review & editing (support). S. Stephan: Conceptualization, Methodology, Supervision, Funding acquisition, Software (support), Writing – review & editing (lead).